OpenStudy (anonymous):

Solve the equation (x-1/x)^2+3(x+1/x)=0 . . . ANSWER = {1±√(3i)/2,-2±√(3)}

OpenStudy (anonymous):

wait i am explaining it all just wait

OpenStudy (anonymous):

\[(x-\frac{ 1 }{ x })^2+3(x +\frac{ 1 }{ x })=0\]

OpenStudy (anonymous):

this is the question

OpenStudy (anonymous):

and this is the answer

hartnn (hartnn):

write (x-1/x)^2 as (x+1/x)^2 -4 and put y= x+1/x

OpenStudy (anonymous):

\[\frac{ 1±\sqrt{3i}}{ 2 },-2±\sqrt{3}\]

OpenStudy (anonymous):

this should e the answer

OpenStudy (anonymous):

do you want us to show the steps?

OpenStudy (anonymous):

nope just want the ideason how to solve it

hartnn (hartnn):

write (x-1/x)^2 as (x+1/x)^2 -4 and put y= x+1/x could u take i from here ? u get a quadratic in y

OpenStudy (anonymous):

its your choice whether u show the steps or not

hartnn (hartnn):

y^2+3y-4=0 solve this

OpenStudy (anonymous):

I think @hartnn gave you the best way.

OpenStudy (anonymous):

\[(\frac{x^2-1}{x^2})^2+3(\frac{x^2+1}{x^2})\]

OpenStudy (anonymous):

ok i am solving the quadratic

OpenStudy (anonymous):

@MrMoose no squares inside brackets of the question

OpenStudy (anonymous):

\[\frac{(x^2-1)^2+3(x^2-1)^2}{x^4}\]

OpenStudy (anonymous):

\[u = x^2\] \[(u-1)^2+3(u+1)^2 = 0\]

OpenStudy (anonymous):

\[4u^2+4u+4=0\]

OpenStudy (anonymous):

\[4(x^4+x^2+1) = 0\]

OpenStudy (anonymous):

\[x^4+x^2+1\] obviously has the factors w and w^2, where they are the 3rd roots of unity

OpenStudy (anonymous):

@hartnn i've solved the quadratic and its coming y=1 and y=-4 whats next

OpenStudy (anonymous):

@MrMoose there is no x^2 in the question how did u get x^2

hartnn (hartnn):

now x+1/x = 1 or x+1/x = -4 so u get x^2+1=x or x^2+1 = -4x now solve these 2 quadratics. preferably using formula

OpenStudy (anonymous):

by getting a common denominator \[(x-\frac{1}{x}) = (\frac{x^2-1}{x}) \]

OpenStudy (anonymous):

the denominator u=is also x^2 as u mentioned above @MrMoose

OpenStudy (anonymous):

the denominator doesn't matter

OpenStudy (anonymous):

if the numerator is 0, the whole thing is 0

OpenStudy (anonymous):

@hartnn the first quadratic solved and it is matching the first answer now checking second one

OpenStudy (anonymous):

done thanks @hartnn

hartnn (hartnn):

welcomes :)

OpenStudy (anonymous):

@hartnn goku charging his spirit bomb?

hartnn (hartnn):

yeah ! give him your energy :P

OpenStudy (anonymous):

given :)